OpenStudy (anonymous):

Reduce (1-a^2)/(a-1) to lowest terms. (a) 1+a (b)-(a-1) (c) -1-a (d) a-1) (e) none of the above

OpenStudy (anonymous):

1-a^2 = (1-a)(1-a) so (1-a^2)/a-1 = (1-a)(1-a)/(a-1) = -(1-a) = a-1 so the answer is E